OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8.

OpenStudy (anonymous):

y = 1/32 ^x2 y^2 = 8x y^2 = 32x y = 1/8 ^x2

OpenStudy (anonymous):

@amistre64 Could you help me with this?

OpenStudy (amistre64):

your notations seem a little off in the options

OpenStudy (amistre64):

but the mechanics involved is such that the definition of a parabola is all the points (x,y) that are the same distance from the focus (a,b) and some line directix ... in this case y=k the distance formula is therefore a good formula to employ if need be: df^2 = (x-a)^2 + (y-b)^2 and dd^2 = (x-x)^2 + (y-k)^2 equating them we get: df^2 = dd^2 (x-a)^2 + (y-b)^2 = (x-x)^2 + (y-k)^2 and simplify with algebra (x-a)^2 + (y-b)^2 = (y-k)^2 (x-a)^2 + y^2 -2by +b^2 = y^2 -2k y +k^2 (x-a)^2 -2by +b^2 = -2k y +k^2 (x-a)^2 +b^2 - k^2 = 2by -2k y (x-a)^2 +b^2 - k^2 = (2b -2k) y (x-a)^2 +b^2 - k^2 ------------------ = y 2 (b-k)

OpenStudy (amistre64):

lets see if the values make this prettier lol

OpenStudy (amistre64):

(a,b) = (0,8); k=-8 (x-0)^2 +8^2 - 8^2 ------------------ = y 2 (8+8) x^2 ------ = y 2 (16)

OpenStudy (anonymous):

Sorry really slow internet, sigh. So it would be the third answer choice? (y^2= 32x)

OpenStudy (anonymous):

Wait no it's y = 1/32 x2?

OpenStudy (amistre64):

the site is definitely not working well at the moment, but yes: y = 1/32 x^2 fits

OpenStudy (anonymous):

I thought it only happened to me, lol. THANKS SO MUCH!